# D-tensor paraproducts and its caricatures

### v1

@@ 4.2 Cone singularity in complex plane (d=3)
$f(z)=\begin{cases}(1-|z|)^{\alpha}&\text{if }|z|<1\\ (1-\frac{1}{|z|})^{\alpha}&\text{if }\frac{1}{|z|}<1\end{cases}$ (78)
@@
$A_{3}(f)=sin((\frac{j_{x}+j_{iy}}{2}<f(z),\psi^{j_{x},j_{iy}^{\prime}}(x,iy)>\psi^{j_{x},j_{iy}^{\prime}}(x,iy))$ (79)
Figure 4: Cone singularity in complex plane: (a - c) $f,A(f),\tilde{A}(f)$ at t=15 (d- f) $f,A(f),\tilde{A}(f)$ at t=70, (g-i) $f,A(f),\tilde{A}(f)$ at t=115.
@@
We process all the figures in Figure 4 with the tensor haar basis. The red squares are the support of the tensor haar functions with the largest coefficients. This exemplifies the ability of the decomposition to recover the singularity in the presence of smooth non-linear distortions - the singularity is detected as expected for the original function, $f$, goes undetected for the non-linear distortion in the middle column, $A(f)$, and is recovered again by the approximation, $\tilde{A}(f)$ in the right column.
5 Acknowledgments
The author wishes to thank Ronald R. Coifman for useful discussions.

### v2

@@ 6. Cone singularity in complex plane (d=3)
(64) $f(z)=\begin{cases}(1-|z|)^{\alpha}&\text{if }|z|<1\\ (1-\frac{1}{|z|})^{\alpha}&\text{if }\frac{1}{|z|}<1\end{cases}$
@@
(65) $A_{3}(f)=sin((\frac{j_{x}+j_{iy}}{2}<f(z),\psi^{j_{x},j_{iy}^{\prime}}(x,iy)>\psi^{j_{x},j_{iy}^{\prime}}(x,iy))$
Figure 1. Cone singularity in complex plane: (a - c) $f,A(f),\tilde{A}_{(N_{i})_{i=0}^{d}}(f)$ at t=15 (d- f) $f,A(f),\tilde{A}_{(N_{i})_{i=0}^{d}}(f)$ at t=70, (g-i) $f,A(f),\tilde{A}_{(N_{i})_{i=0}^{d}}(f)$ at t=115.
@@
We process all the figures in Figure 1 with the tensor haar basis. The red squares are the support of the tensor haar functions with the largest coefficients. This exemplifies the ability of the decomposition to recover the singularity in the presence of smooth non-linear distortions - the singularity is detected as expected for the original function, $f$, goes undetected for the non-linear distortion in the middle column, $A(f)$, and is recovered again by the approximation, $\tilde{A}_{(N_{i})_{i=0}^{d}}(f)$ in the right column.